OpenStudy (anonymous):

Lim (9-x)/(3-√x), x ->9. Correct answer is 6

OpenStudy (princeharryyy):

lim ((3- sqrt(x))*(3+ sqrt(x)))/(3- sqrt(x)) =>lim (3- sqrt(x)) plugiin' in the value of x we get 3+sqrt(9) =>3+3 =>6 answer @Gloria63 where sqrt is equal to square root

OpenStudy (anonymous):

\(\displaystyle \lim_{x\rightarrow 9}~\frac{9-x}{3-\sqrt{x}}\) using LHS \(\displaystyle \lim_{x\rightarrow 9}~\frac{1}{\frac{1}{2\sqrt{x}}}\) \(\displaystyle \lim_{x\rightarrow 9}~2\sqrt{x}\)

OpenStudy (anonymous):

i cancelled negatives if anything

OpenStudy (anonymous):

should suffice:)

OpenStudy (anonymous):

Tyvm to both of you. First answer is really easy

OpenStudy (anonymous):

LHS goes quite handy, but if you prefer rationalizing then y not:)